Moon Gazers, Inc., produces telescopes for use by high school students. All direct materials used in the production of telescopes are added at the beginning of the manufacturing process. Labor and overhead are added evenly thereafter, as each unit is assembled, adjusted, and tested. Moon Gazers uses process costing and had the following unit production information available for the months of January and February. Jan. Feb.

Jan. Feb.

Number of units in beginning work in process inventory 0 100

Number of units started during the month 400 500

Total number of units transferred to finished goods 300 140

The units remaining in work in process at the end of January were approximately 60 percent complete. During the month of February, all of the beginning work in process units were completed and the units remaining in work in process at the end of the month were approximately 80 percent complete.

a. For the month of January, calculate the equivalent units produced for each of the two cost categories—direct materials and labor and overhead.

b. For the month of February, calculate the equivalent units produced for each of the two cost categories—direct materials and labor and overhead.

Equivalent Units Produced January    February
a. Direct materials
b. Labor and overhead
Answer #1

calculate equivalent unit

Equivalent unit = Unit completed+(Ending work in process*Percent completion)

Equivalent units produced January February
a Direct material 300+100 = 400 140+460 = 600
b Labor and overhead 300+(100*60%) = 360 140+(460*80%) = 508